# Demonstration of Hardware Efficient Photonic Variational Quantum Algorithm

### v1

@@ Abstract
Quantum computing has brought a paradigm change in computer science, where non-classical technologies have promised to outperform their classical counterpart. Such an advantage was only demonstrated for tasks without practical applications, still out of reach for the state-of-art quantum technologies. In this context, a promising strategy to find practical use of quantum computers is to exploit hybrid quantum-classical models, where a quantum device estimates a hard-to-compute quantity, while a classical optimizer trains the parameters of the model. In this work, we demonstrate that single photons and linear optical networks are sufficient for implementing Variational Quantum Algorithms, when the problem specification, or ansatz, is tailored to this specific platform. We show this by a proof-of-principle demonstration of a variational approach to tackle an instance of a factorization task, whose solution is encoded in the ground state of a suitable Hamiltonian. This work which combines Variational Quantum Algorithms with hardware efficient ansatzes for linear-optics networks showcases a promising pathway towards practical applications for photonic quantum platforms.
@@ I Introduction
In the past decade, quantum computing has risen to the forefront of current technological discussions, fundamentally transforming the field of computation. Indeed, quantum technologies promise to address computational problems exponentially faster and more efficiently, when being compared to their classical counterparts, both for research scopes, e.g. simulating complex physical systems [1, 2], as well as for tackling real-world problems [3, 4]. In this context, two specific quantum computing paradigms that have garnered popularity in recent years are adiabatic quantum computing (AQC) and Variational Quantum Algorithms (VQAs) [5, 6]. The first draws inspiration from the concept of adiabatic evolution and it maps the problem of interest onto finding the ground state of a given Hamiltonian ($H_{f}$). Since such a state can be experimentally very complex to achieve, the algorithm starts from the implementation of another state (easier to prepare), which corresponds to the ground state of an initial Hamiltonian $H_{i}$. Then, the latter is slowly (and in a controlled way) modified, so that the physical system is maintained in the ground state. In the end, when $H_{i}$ converges to $H_{f}$, it is possible to retrieve the solution of the considered computational problem. Despite its robust nature, obtaining fruitful outcomes using AQC on noisy intermediate scale quantum devices (NISQ) is a challenging task, as adiabatically modifying the initial Hamiltonian requires long sequences of quantum gates [7].
@@
Here, we implement a VQA, which is custom-tailored to photonic platforms, to address a proof-of-principle instance of the factorization problem. Let us note that our method for factorization, analogously to the one based on AQC, has a better scalability, in terms of the number of required qubits, when being compared to Shor’s algorithm [25]. In detail, we employ four input/output ports of a universal photonic integrated circuit (PIC) consisting of a mesh of programmable interferometers that can implement arbitrary unitary transformations on Fock input states. This kind of platforms is a natural candidate for adaptive computational algorithms, benefitting from their ease of encoding and manipulating information with high fidelity. Here we present such a protocol where we map our problem to the search of the ground state of a suitable Hamiltonian and implement a gradient based optimization on our photonic platform to obtain the experimental parameters yielding the correct solution. A schematic of this process is depicted in Figure 1.
@@ II Results
where $\mathcal{F}=\lceil\sqrt{N}\rceil$ (note that $\lceil a\rceil$ indicates the largest integer smaller or equal to $a$) and $m(b)$ denotes the smallest number of bits required for representing $b$. The offset $\delta N$, instead, is required because the gap between two successive primer numbers is not constant and does not grow linearly with $N$ and it amounts to $\sqrt{\mathcal{F}^{2}-N}$ (see Supplemental Material section I). This implies that the solution can be expressed in the following as a binary number (bit-string):
$x_{c}=x_{n_{x}-1}\dots x_{1}x_{0},\;\;\;\;\;\;\;\;y_{c}=y_{n_{x}-1}\dots y_{1}y_{0}.$ (3)
However, when the number $N$ to factorize is given by the product of two prime numbers, the number of unknown bits is lower, if we consider that each odd binary number ends with $1$. Therefore, $n_{x}$ ($n_{y}$) is reduced to $n_{x}-1$ ($n_{y}-1$) and, if we want to encode $x_{c}$ and $y_{c}$ in a quantum state, the required Hilbert space dimension results $2^{n_{x}+n_{y}-2}$.
In our implementation, anyway, for practical purposes we can reduce the number of unknown bits even further. Indeed, if we consider the factorization of the number $35=5\times 7$, the length of the bit-strings $x$ and $y$ constituting the prime factors must amount to $n_{x}=n_{y}=3$ (as detailed in [21] and as reported in Supplemental Material section I). This implies that also the most significant bit of the factors must be 1, otherwise a lower number of bits would be sufficient. So, in this specific case, the total number of unknown bits is further reduced, because, our factors $x_{c}$ and $y_{c}$ will have the form $11x_{1}$ 1 and $11y_{1}$ 1 respectively. Let us note that this restriction holds only in our case, but the protocol can handle more general cases. Indeed, it is possible to evaluate the maximum number of required bits (as reported in Supplemental Material section I), which will be the number of bits outputted by the method. Hence, if a lower amount of bits is sufficient, the most significant ones will be 0. In the present scenario, the Hamiltonian that encodes the problem can be written in an Ising-like form as follows:
@@
where $N$ is the number to factorize, $I$ is the identity operator, $\Pi_{z}^{1}$ refers to the projector onto the eigenvector of the $\sigma_{Z}$ operator corresponding to outcome $1$ and the subsacript indicates the qubit on which it acts (see the Supplemental Material section II for further details). Then, the state that will minimize Eq. (4) will be represented by a two-qubit state |$\ket{x_{1}y_{1}}$, where the two unknown bits will be $x_{1}$ and $y_{1}$. This is due to the form of the Hamiltonian, because the two factors $x$, $y$ are expressed, in binary form, as follows: $14+2x_{1}+$ 1 and $14+2y_{1}+$ 1. So, the ground state of $H_{p}$ will be the one for which the expected values of the projector onto the eigenstate $|1\rangle$ of the operator $\sigma_{z}$ amount to $x_{1}$ and $y_{1}$.
@@
Such a single photon input is generated by type-II Spontaneous parametric down conversion (SPDC) in a periodically poled $\mathrm{KTiOPO_{4}}$-crystal in a single-pass configuration as depicted in Figure 3. Then, one photon is directly sent to a detection unit consisting of Superconducting Nanowire Single Photon Detectors (SNSPDs). It thereby heralds the presence of its twin, which is routed to a tunable optical circuit. This is the core of our experimental apparatus and it is constituted by a six-mode universal PIC, featuring only linear optical elements, i.e. phase shifters and beam splitters, whose architecture follows the optimal design proposed in [8]. In detail, the circuit consists of a grid of variable beam splitters, i.e. Mach-Zehnder interferometers (see Figure 2), each characterized by a transformation matrix denoted as $T_{jk}(\theta,\phi)$, where $\theta$ determines the reflectivity and $\phi$ is an external phase. With a specific ordered sequence $S$ of these transformation matrices, it is possible to create a programmable unitary operator described as:
$U(\vartheta,\varphi)=D\prod_{j,k\in S,l=\{1,2,\dots L\}}T_{jk}(\theta_{l},\phi_{l})$ (6)
@@
The PIC is manufactured in a borosilicate glass substrate through femtosecond laser waveguide writing [35]. The tunability is achieved by fabricating thermo-optic phase shifters (i.e. resistive heaters) in a thin film of gold deposited on the chip surface. The device exhibits an overall transmission of $54\%$ and can implement arbitrary unitary operations on six modes with an average amplitude fidelity of 0.9970 ± 0.0017 [36].
For our experiment, we need to use only four input and output modes, that are tunable by acting on twelve phase shifters (L=6), while the surrounding actuators are set in reflection ($\theta=\pi/2$) to isolate these modes from the others available in the PIC. Our ansatz is custom-tailored to these structures. Hence, to find these two sets of optimal parameters $(\vartheta_{1},\varphi_{1})$ and $(\vartheta_{2},\varphi_{2})$, we perform $117$ repetitions of the experiment, starting with random initial parameters. For each repetition, we initialize our ansatz by applying twelve random phase shifts ($L=6$) and injecting the input state as $|\psi_{0}\rangle=|00\rangle$, namely a single photon in the first mode. Then, we collect the photon statistics at each output, corresponding to a projection of the rotated input onto the computational basis, comprising the states $|00\rangle$, $|01\rangle$, $|10\rangle$ and $|11\rangle$. Following this, we post-process the data to evaluate the energy of |$\ket{\psi_{0}}$ with respect to $H_{P}$ on a classical computer, using Eq. (5). Note that, for $H_{p}$, which only involves expected values in the $\sigma_{z}$ basis, a full tomography on the output state is not required to achieve the correct value of Eq. (5) (see the Supplemental Material section III).
@@
It is worthwhile to point out that the classical optimization algorithm, in principle, requires no knowledge about $U(\vartheta,\varphi)$ or even the possibly nonlinear mapping between the phases and their control currents. In our case, though, we exploited a pre-characterization of the device, to ensure that no light would go to the undetected modes of the 6 input/ 6 output circuit.
@@
The convergence of the energy value with respect to the function evaluations is depicted in Figure 4, where the different curves correspond to different initial phase configurations and the dashed line represents the theoretical limit. It can be clearly seen that the optimization works smoothly and is able to reach the the ground state in every instance with an accuracy ranging from $99\%$ to $99.99\%$. Furthermore, let us point out that the ground state of $H_{p}$ is two-fold degenerate and each instance of the experiment converges to one particular gound state corresponding to one optimal unitary i.e., either $U(\vartheta_{1},\varphi_{1})$ or $U(\vartheta_{2},\varphi_{2})$This is visible throughout the 117 repetitions of the experiment, whose averaged probabilities are represented in Figure 5a (purple bars), where we also compare them to the theoretical predictions (in blue). We obtained the solutions |$\ket{01}$ and |$\ket{10}$ with $50.7\%$ and $44.6\%$ respectively. This implies, as previously mentioned, that the unknown bits for our two factors are |$\ket{x_{1}y_{1}}=\ket{10}$ or, equivalently, |$\ket{01}$.
@@
Figure 5: Experimental factorization results. a) The bars correspond to the experimental frequencies with which a photon is detected at each of the four outputs of our circuit. The reported values are obtained by averaging over the results of $117$ different experimental optimizations, starting from different configurations. For our encoding, injecting one photon in the $i$-th mode constitutes the $i$-th element of a 2-qubits computational basis. Therefore, the bars representing the highest frequencies indicate the ground state |$\ket{x_{1}y_{1}}$ of the Hamiltonian outlined in Eq. (4), which is an almost equal superposition of |$\ket{01}$ and |$\ket{10}$. This effectively resolves the factorization problem, allowing us to construct the factors as $51x_{1}1=101=$ 5 and $71y_{1}1=111=$ 7, or vice versa. b) As all states of the form $\sqrt{\alpha}\ket{01}+\sqrt{1-\alpha}\ket{10}$ exhibit the same energy with respect to $H_{p}$ (see supplemental material Sec. 3), every individual optimization converges to one arbitrary of those possible outcomes. The average output of independent repetitions thus approaches the equal superposition $\alpha=0.5$, as can be seen on the fidelity of $0.97546(5)$ towards this state after 117 repetitions (violet), while the fidelities to both |$\ket{01}$ (green) and |$\ket{10}$ (teal) concordantly revolve around $0.7$. The fidelity (Eq. (7)) is evaluated by interpreting the output distribution measured in the computational basis as normalized state vector. For both plots, the uncertainties (shaded regions for 2 standard deviations), evaluated considering an underlying Poissonian statistics, using a Monte Carlo Simulation and are too small to be visible.
@@ III Discussion
This work reports on the implementation of a VQA on an integrated photonic circuit, exploiting an ansatz which is tailored to our platform. In particular, we present a proof-of-principle demonstration for the factorization of $35=5\times 7$, executed on a four-input/output modes of a universal PIC. Our strategy involves the use of a qudit with four levels, corresponding to the four modes of the photonic circuit. This is then interpreted as two qubits, encoding the solution in the ground state of an Ising Hamiltonian. Subsequently, we determine its energy through a hybrid quantum-classical optimization problem, utilizing gradient descent. In further detail, we design an ansatz specifically for state-of-the-art PICs that allow high-quality processing of single photons by using only passive optical elements [8]. Our encoding of the two-qubit computational basis involves the binary representation of the mode from which one single photon enters an optical circuit, where the first mode corresponds to $|00\rangle$, the second to $|01\rangle$, and so forth. At each step of the optimization, we collect the statistics obtained by projecting the evolved input state onto the computational basis and post-process it to retrieve the energy value. Then, we identify a new configuration of our circuit and we iterate the procedure, until it converges to the minimum energy value, corresponding to the ground state of our Hamiltonian. This allows us to reach the correct solution for our factorization problem with a success probability exceeding 98%, in fewer than 150 iterations and for $117$ different starting conditions, highlighting its reduced sensitivity to initial parameters. This study marks the first instance of factorization on an optical computing device and shows the utility of the photonic devices in Variational Quantum Algorithms, showcasing advantages in terms of resource requirements and resilience against inherent noises. Although our demonstration is dedicated to the factorization problem, we want to emphasize that our approach can readily be employed to address other challenges, given that our results indicate that universal integrated circuits are a suitable platform for the implementation of VQAs. This is significant in the landscape of quantum technologies, for their versatility, high achievable fidelity in the implementation of arbitrary unitary transformations and scalability. These features make photonics particularly apt for adaptive protocols, e.g. the training of machine learning algorithms. On the other hand, the limitations of this platform are related to the difficulty of implementing two-qubit gates and hence non-linear operations. Indeed, the fact that photons tend not to interact with the environment implies that two-qubit gates can only be implemented probabilistically and heralded through ancillary photons or post-selection [37]. These issues were circumvented in our case by encoding $n$ qubits in $2^{n}$ modes, offering a platform which is exactly equivalent to the quantum circuit formalism. The drawback of this approach amounts to the number of spatial modes scaling exponentially with the dimension of the system. However, this can potentially be eased by future developments, as suggested by several recent studies investigating the possibility of implementing non-linear operations through encoding, feedback loops and enhancing the expressivity of the circuit through quantum interference [38, 39, 40]. Moreover, recent studies have highlighted that optical computational models could grant lower energy consumption, with respect to standard models [41].

### v2

@@ Abstract
Quantum computing has changed the paradigm of computer science, where quantum technologies have promised to outperform classical ones. Such an advantage was only demonstrated for tasks with no application or out of reach for the state-of-art quantum technologies. In this context, a promising strategy to find practical use of quantum computers exploits hybrid models, where a quantum device estimates a hard-to-compute quantity, while a classical optimizer trains the parameters. In this work, we demonstrate that single photons and linear optical networks are sufficient for implementing Variational Quantum Algorithms, when the problem specification (ansatz), is tailored to this specific platform. We show this by a proof-of-principle demonstration to tackle an instance of a factorization task, whose solution is encoded in the ground state of a suitable Hamiltonian. This work which combines Variational Quantum Algorithms with hardware efficient ansatzes for linear-optics showcases a promising pathway towards practical applications for photonic quantum platforms.
@@ I Introduction
In the past decade, quantum computing has risen to the forefront of current technological discussions, fundamentally transforming the field of computation.
@@
Indeed, quantum technologies promise to address computational problems exponentially faster and more efficiently, when being compared to their classical counterparts, both for research scopes, e.g. simulating complex physical systems [1, 2], as well as for tackling real-world problems [3, 4]. In this context, two specific quantum computing paradigms that have garnered popularity in recent years are adiabatic quantum computing (AQC) and Variational Quantum Algorithms (VQAs) [5, 6]. The first draws inspiration from the concept of adiabatic evolution and it maps the problem of interest onto finding the ground state of a given Hamiltonian ($H_{f}$). Since such a state can be experimentally very complex to achieve, the algorithm starts from the implementation of another state (easier to prepare), which corresponds to the ground state of an initial Hamiltonian $H_{i}$. Then, the latter is slowly (and in a controlled way) modified, so that the physical system is maintained in the ground state. In the end, when $H_{i}$ converges to $H_{f}$, it is possible to retrieve the solution of the considered computational problem. Despite its robust nature, obtaining fruitful outcomes using AQC on noisy intermediate scale quantum devices (NISQ) is a challenging task, as adiabatically modifying the initial Hamiltonian requires long sequences of quantum gates [7].
@@
Here, we implement a VQA, which is custom-tailored to photonic platforms, to address a proof-of-principle instance of the factorization problem. Regarding the adopted platform, we employ four input/output ports of a universal photonic integrated circuit (PIC) consisting of a mesh of programmable interferometers that can implement arbitrary unitary transformations on Fock input states. This kind of platforms is a natural candidate for adaptive computational algorithms, benefitting from their ease of encoding and manipulating information with high fidelity. Here we present such a protocol where we map our problem to the search of the ground state of a suitable Hamiltonian and implement a gradient based optimization on our photonic platform to obtain the experimental parameters yielding the correct solution. A schematic of this process is depicted in Figure 1. Albeit choosing it as a benchmark, our approach’s scope goes beyond that task. Yet, our method for factorization has a better scalability, in terms of the number of required qubits, when being compared to Shor’s algorithm and, in the special case of a number whose factors are successive prime numbers, also compared to the adiabatic quantum factoring proposed in [25] (See Supplementary Material Section III). As such, our method is in particular applicable on noisy small and medium-sized quantum hardware, which are within the reach of current technologies.
@@ II Results
where $\mathcal{F}=\lfloor\sqrt{N}\rfloor$ (with $\lfloor a\rfloor$ indicating the largest integer smaller or equal to $a$) and $m(b)$ denotes the smallest number of bits required for representing $b$. The offset $\delta N$, instead, is required because the gap between two successive primer numbers is not constant and does not grow linearly with $N$ [35] and it amounts to $\sqrt{\mathcal{F}^{2}-N}$ (see Supplemental Material section I). This implies that the solution can be expressed in the following as a binary number with the maximal bit-string length upper bounded by $n_{x}$, for the first factor, and $n_{y}$ for the second:
$x_{c}=x_{n_{x}-1}\dots x_{1}x_{0},\;\;\;\;\;\;\;\;y_{c}=y_{n_{y}}-1\dots y_{1}y_{0}.$ (3)
However, when the number $N$ to factorize is given by the product of two prime numbers, the number of unknown bits is lower, if we consider that each odd binary number ends with $1$. Therefore, $n_{x}$ ($n_{y}$) is reduced to $n_{x}-1$ ($n_{y}-1$) and, if we want to encode $x_{c}$ and $y_{c}$ in a quantum state, the required Hilbert space dimension results $2^{n_{x}+n_{y}-2}$. Also in this case, these bit-lengths are an overestimation of the real length of the factors.
In our implementation, anyway, for practical purposes we can reduce the number of unknown bits even further. Indeed, if we consider the factorization of the number $35=5\times 7$, which is equal to the largest number factorized through a simplified version of Shor’s algorithm [36], which exploited 7 qubits and succeeded roughly $14\%$ of the cases, the length of the bit-strings $x$ and $y$ constituting the prime factors must amount to $n_{x}=n_{y}=3$ (as detailed in [21] and as reported in Supplemental Material section I). This implies that also the most significant bit of the factors must be 1, otherwise a lower number of bits would be sufficient. So, in this specific case, the total number of unknown bits is further reduced, because, our factors $x_{c}$ and $y_{c}$ will have the form $11x_{1}$ 1 and $11y_{1}$ 1 respectively. Let us note that this restriction holds only in our case, but the protocol can handle more general cases. Indeed, it is possible to evaluate the maximum number of required bits (as reported in Supplemental Material section I), which will be the number of bits outputted by the method. Hence, if a lower amount of bits is sufficient, the most significant ones will be 0. In the present scenario, the Hamiltonian that encodes the problem can be written in an Ising-like form as follows:
@@
where $N$ is the number to factorize, $I$ is the identity operator, $\Pi_{z}^{1}$ refers to the projector onto the eigenvector of the $\sigma_{Z}$ operator corresponding to outcome $1$ and the subscript indicates the qubit on which it acts (see the Supplemental Material section II for further details). Then, the state that will minimize Eq. (4) will be represented by a two-qubit state |$\ket{x_{1}y_{1}}$, where the two unknown bits will be $x_{1}$ and $y_{1}$. This is due to the form of the Hamiltonian, because the two factors $x$, $y$ are expressed, in binary form, as follows: $14+2x_{1}+$ 1 and $14+2y_{1}+$ 1. So, the ground state of $H_{p}$ will be the one for which the expected values of the projector onto the eigenstate $|1\rangle$ of the operator $\sigma_{z}$ amount to $x_{1}$ and $y_{1}$.
@@
[…]
$U(\vartheta,\varphi)=D\prod_{j,k\in S,l}T_{jk}(\theta_{l},\phi_{l});\quad l=\{1,2,\dots L\},$ (6)
@@
The PIC is manufactured in a borosilicate glass substrate through femtosecond laser waveguide writing [37]. The tunability is achieved by fabricating thermo-optic phase shifters (i.e. resistive heaters) in a thin film of gold deposited on the chip surface. The device exhibits an overall transmission of $54\%$ and can implement arbitrary unitary operations on six modes with an average amplitude fidelity of 0.9970 ± 0.0017 [38]. While the employed PIC has a negligible polarization extinction ratio (PER), the path encoding scheme solely relies on a homogenious PER along the PICs width to maintain interference.
For our experiment, we need to use only four input and output modes, that are tunable by acting on twelve phase shifters (L=6), while the surrounding actuators are set in reflection ($\theta=\pi/2$) to isolate these modes from the others available in the PIC. Our ansatz is custom-tailored to these structures. Hence, to find these two sets of optimal parameters $(\vartheta_{1},\varphi_{1})$ and $(\vartheta_{2},\varphi_{2})$, we perform $117$ repetitions of the experiment, starting with random initial parameters. For each repetition, we initialize our ansatz by applying twelve random phase shifts ($L=6$) and injecting the input state as $|\psi_{0}\rangle=|00\rangle$, namely a single photon in the first mode. Then, we collect the photon statistics at each output, corresponding to a projection of the rotated input onto the computational basis, comprising the states $|00\rangle$, $|01\rangle$, $|10\rangle$ and $|11\rangle$. Following this, we post-process the data to evaluate the energy of |$\ket{\psi_{0}}$ with respect to $H_{P}$ on a classical computer, using Eq. (5). Note that, for $H_{p}$, which only involves expected values in the $\sigma_{z}$ basis, a full tomography on the output state is not required to achieve the correct value of Eq. (5) (see the Supplemental Material section IV).
@@
[…]
@@
The convergence of the energy value with respect to the function evaluations is depicted in Figure 4, where the different curves correspond to different initial phase configurations and the dashed line represents the theoretical limit. It can be clearly seen that the optimization works smoothly and is able to reach the ground state in every instance with an accuracy ranging from $99\%$ to $99.99\%$. Furthermore, let us point out that the ground state of $H_{p}$ is two-fold degenerate and each instance of the experiment converges to one particular ground state corresponding to one optimal unitary i.e., either $U(\vartheta_{1},\varphi_{1})$ or $U(\vartheta_{2},\varphi_{2})$This is visible throughout the 117 repetitions of the experiment, whose averaged probabilities are represented in Figure 5a (purple bars), where we also compare them to the theoretical predictions (in blue). We obtained the solutions |$\ket{01}$ and |$\ket{10}$ with $50.7\%$ and $44.6\%$ respectively. This implies, as previously mentioned, that the unknown bits for our two factors are |$\ket{x_{1}y_{1}}=\ket{10}$ or, equivalently, |$\ket{01}$.
@@
Figure 5: Experimental factorization results. a) The bars correspond to the experimental frequencies with which a photon is detected at each of the four outputs of our circuit. The reported values are obtained by averaging over the results of $117$ different experimental optimizations, starting from different configurations. For our encoding, injecting one photon in the $i$-th mode constitutes the $i$-th element of a 2-qubits computational basis. Therefore, the bars representing the highest frequencies indicate the ground state |$\ket{x_{1}y_{1}}$ of the Hamiltonian outlined in Eq. (4), which is an almost equal superposition of |$\ket{01}$ and |$\ket{10}$. This effectively resolves the factorization problem, allowing us to construct the factors as $51x_{1}1=101=$ 5 and $71y_{1}1=111=$ 7, or vice versa. b) As all states of the form $\sqrt{\alpha}\ket{01}+\sqrt{1-\alpha}\ket{10}$ exhibit the same energy with respect to $H_{p}$ (see Supplemental Material section IV), every individual optimization converges to one arbitrary of those possible outcomes. The average output of independent repetitions thus approaches the equal superposition $\alpha=0.5$, as can be seen on the fidelity of $0.97546(5)$ towards this state after 117 repetitions (violet), while the fidelities to both |$\ket{01}$ (green) and |$\ket{10}$ (teal) concordantly revolve around $0.7$. The fidelity (Eq. (7)) is evaluated by interpreting the output distribution measured in the computational basis as normalized state vector. For both plots, the uncertainties (shaded regions for 2 standard deviations), evaluated considering an underlying Poissonian statistics, using a Monte Carlo Simulation and are too small to be visible.
@@ III Discussion
This work reports on the implementation of a VQA on an integrated photonic circuit, exploiting an ansatz which is tailored to our platform. In particular, we present a proof-of-principle demonstration for the factorization of $35=5\times 7$, executed on a four-input/output modes of a universal PIC. Our strategy involves the use of a qudit with four levels, corresponding to the four modes of the photonic circuit. This is then interpreted as two qubits, encoding the solution in the ground state of an Ising Hamiltonian. Subsequently, we determine its energy through a hybrid quantum-classical optimization problem, utilizing gradient descent. In further detail, we design an ansatz specifically for state-of-the-art PICs that allow high-quality processing of single photons by using only passive optical elements [8]. Our encoding of the two-qubit computational basis involves the binary representation of the mode from which one single photon enters an optical circuit, where the first mode corresponds to $|00\rangle$, the second to $|01\rangle$, and so forth. At each step of the optimization, we collect the statistics obtained by projecting the evolved input state onto the computational basis and post-process it to retrieve the energy value. Then, we identify a new configuration of our circuit and we iterate the procedure, until it converges to the minimum energy value, corresponding to the ground state of our Hamiltonian. This allows us to reach the correct solution for our factorization problem with a success probability exceeding 98%, in fewer than 150 iterations and for $117$ different starting conditions, highlighting its reduced sensitivity to initial parameters. This study marks the first instance of factorization on an optical computing device and shows the utility of the photonic devices in Variational Quantum Algorithms, showcasing advantages in terms of resource requirements and resilience against inherent noises. This is reflected by the fact that, with the integrated platforms currently available on the market, our implementation could already factorize a semi-prime number up to $93$, while considering the product of two consecutive primes, this limit goes up to $899$. These numbers are well beyond what has current been achieved with Shor’s algorithm [36] and this limit is likely to increase with recent developments in integrated photonics. Indeed, until now, the largest number factorized through a simplified version of Shor’s algorithm is $35$, by exploiting 7 qubits and obtained a success rate of roughly $14\%$ of the cases [36]. The number of required qubits was then lowered down in a more recent work, where $21$ was factored using 5 qubits [39]. Moreover, although our demonstration is dedicated to the factorization problem, we want to emphasize that our approach can readily be employed to address other challenges, given that our results indicate that universal integrated circuits are a suitable platform for the implementation of VQAs. Let us note that the integrated photonic platforms that are currently on the market are already sufficient to encode up to 5 qubits using path-encoded single-photon architectures. This places our hardware-efficient variational framework well within the range of many established benchmark problems in near-term quantum computing. For instance, ground-state estimation of simple molecular systems such as $\mathrm{H_{2}}$ using VQE requires only few qubits [40, 41], and combinatorial problems like MaxCut on 3 to 4 node graphs can also be encoded on systems smaller than 5 qubits [42]. These examples are representative of diverse problem classes and underscore that our method is not inherently limited to factorization tasks, but is broadly applicable within the capabilities of current photonic hardware. Moreover, the versatility, high achievable fidelity in the implementation of arbitrary unitary transformations and scalability of photonics platforms make them particularly apt for adaptive protocols, e.g. the training of machine learning algorithms. On the other hand, the limitations of this platform are related to the difficulty of implementing two-qubit gates and hence non-linear operations. Indeed, the fact that photons tend not to interact with the environment implies that two-qubit gates can only be implemented probabilistically and heralded through ancillary photons or post-selection [43]. These issues were circumvented in our case by encoding $n$ qubits in $2^{n}$ modes, offering a platform which is exactly equivalent to the quantum circuit formalism. While this approach offers the advantage of the number of input photons staying constant - specifically one - independent of problem size, it exhibits the drawback of the number of necessary PIC modes scaling exponentially in the length of the encoded bit strings. Consequently, the number of parameters to optimize also increases exponentially. Furthermore, the reliance on a single photon relaxes requirements on PIC transmittivity and detection efficiency. As these quantities affect the integration time exponentially in the number of photons, they only impact our protocol in a constant manner. However, the drawbacks can potentially be eased by future developments, as suggested by several recent studies investigating the possibility of implementing non-linear operations through encoding, feedback loops and enhancing the expressivity of the circuit through quantum interference [44, 45, 46]. Moreover, recent studies have highlighted that optical computational models could grant lower energy consumption, with respect to standard models [47].